# Proactive Model Adaptation Against Concept Drift for Online Time Series Forecasting

### v1

@@ Abstract.
Time series forecasting always faces the challenge of concept drift, where data distributions evolve over time, leading to a decline in forecast model performance. Existing solutions are based on online learning, which continually organize recent time series observations as new training samples and update model parameters according to the forecasting feedback on recent data. However, they overlook a critical issue: obtaining ground-truth future values of each sample should be delayed until after the forecast horizon. This delay creates a temporal gap between the training samples and the test sample. Our empirical analysis reveals that the gap can introduce concept drift, causing forecast models to adapt to outdated concepts. In this paper, we present Proceed, a novel proactive model adaptation framework for online time series forecasting. Proceed first operates by estimating the concept drift between the recently used training samples and the current test sample. It then employs an adaptation generator to efficiently translate the estimated drift into parameter adjustments, proactively adapting the model to the test sample. To enhance the generalization capability of the framework, Proceed is trained on synthetic diverse concept drifts. We conduct extensive experiments on five real-world datasets across various forecast models. The empirical study demonstrates that our proposed Proceed brings more performance improvements than the state-of-the-art online learning methods, significantly facilitating forecast models’ resilience against concept drifts.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2025; Barcelona, Spain.isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data stream miningccs: Computing methodologies Online learning settings
@@ 1. Introduction
Time series forecasting has been a prevalent task in numerous fields such as energy (Boussif et al., 2023), retail (Böse et al., 2017), climate (Bi et al., 2023) and finance (Zhao et al., 2023). The past decade has witnessed a surge of deep learning-based forecast models (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2023) that take past time series observations to predict values in the next $H$ steps, where $H$ is called forecast horizon. Due to the dynamic nature of the environment, latent concepts that influence observation values (e.g., social interest (Bai et al., 2023), stock market sentiment (Li et al., 2022)) often change over time. This ubiquitous phenomenon is known as concept drift (Gama et al., 2014). In the presence of concept drift, future test time series may not follow a similar data distribution as the historical training data, causing degradation in forecasting performance (You et al., 2021; Zhao et al., 2023).
@@
Online learning is commonly used to mitigate the effects of concept drift. The key consideration is to continually transform the newly observed time series into a set of training samples and adjust model parameters by minimizing the forecast errors on the new training samples. In addition to the standard fine-tuning technique, recent works (Pham et al., 2023; Zhang et al., 2023) proposed more advanced model adaptation techniques, which focus on how to effectively adapt to recent data by utilizing forecasting feedback (e.g., forecast errors or gradients) on the new training samples. Among them, FSNet (Pham et al., 2023) monitors the gradients in previous fine-tuning, transforms them into parameter adjustments, and tailors a new forecast model to the current training samples. OneNet (Zhang et al., 2023) is an online ensembling network that generates ensemble weights to combine two forecast models and dynamically adjusts the ensemble weights and the models’ parameter weights according to the forecast errors. However, existing online learning methods overlook the fact that the ground truth of each prediction is not available immediately but is delayed after the forecast horizon.
@@
As illustrated in Fig. 1, an online learning task inherently has a $H$-step feedback delay in time series forecasting, resulting in a temporal gap (at least $H$ steps) between available training samples and the test sample. Formally, at each online time $t$, the current horizon window $\mathbf{Y}_{t}$, which have not been observed yet, has overlap with $\mathbf{Y}_{t-H+1},\dots,\mathbf{Y}_{t-1}$. These samples cannot provide complete forecasting feedback and supervised signals for the performant, prevailing time series forecasting models which predict $H$-step values directly (Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024). Hence, the available training samples for online learning at time $t$ is $\mathcal{D}_{t-}=\{(\mathbf{X}_{t^{\prime}},\mathbf{Y}_{t^{\prime}})\mid t^{\prime}\leq t-H\}$, which has a temporal distance to the test sample $(\mathbf{X}_{t},\mathbf{Y}_{t})$. The temporal gap issue suggests that addressing concept drift by fitting the forecast model to the recent training samples may be insufficient, since concept drift may also occur over the $H$-step temporal gap between the recent training samples and the current test sample.
@@
Figure 1. (a) Example of feedback delay when $H=3$. At online time $t=5$, we use observation values $\mathbf{X}_{5}=\{\mathbf{v}_{1},\cdots,\mathbf{v}_{5}\}$ to forecast future values $\mathbf{Y}_{5}=\{\mathbf{v}_{6},\mathbf{v}_{7},\mathbf{v}_{8}\}$, while the ground truth is known until $t=8$ and the forecasting feedback arrives with 3-step delay. (b) A temporal gap always exists between new training samples $\mathcal{D}_{t-}$ and the test sample $\mathbf{X}_{t}$, where concept drift may occur.
@@
To investigate the impact of this temporal gap on forecasting performance, we examine two online learning strategies using real-world time series datasets (detailed in Section 3). The first strategy fine-tunes model parameters using the latest available training sample at time $t$, which is $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$. The second strategy (Pham et al., 2023; Zhang et al., 2023) omits the feedback delay and adapts the model using $(\mathbf{X}_{t-1},\mathbf{Y}_{t-1})$, which is infeasible in practice due to future information leakage. The empirical results show that the average forecast error of the first strategy is approximately double that of the second strategy, indicating that even the latest observed time series pattern differs from the test sample notably. Moreover, the performance gap between the two strategies becomes more significant as the forecast horizon increases, suggesting that the concept drift becomes more pronounced over longer time intervals. These findings highlight the presence of a substantial concept drift between the practical training samples and the test sample, limiting the effectiveness of current online learning techniques that adapt model parameters to the potentially outdated concepts and leave the concept drift unresolved.
@@
To address these issues, we propose a proactive model adaptation framework named Proceed that responds to the concept drift before forecasting the test sample. In scenarios where optimal model parameters changes with time, we posit that such parameter changes are affected by the drift across a latent concept space, i.e., there may exist learnable relationships between the direction and degree of concept drift and those of parameter changes. In light of this, we propose to infer parameter changes based on concept drift, rather than directly generating a whole model. Specifically, Proceed begins with a forecast model that have fit new training samples by gradient descent. Given a test sample, Proceed exploits latent features that encapsulate the concept drift between the training samples and the test sample. In response to the undergoing concept drift, Proceed generates parameter adjustments via bottleneck layers, which yield a compact set of adaptation coefficients that rescale all model parameters. In this way, our solution eschews direct mapping from concept drift to per-parameter adjustments, instead opting for a more nuanced and efficient strategy. We expect the adapted model with rescaled parameters to approach the optimum for the test sample. To enhance the generalization ability of Proceed, we shuffle historical data to synthesize diverse concept drifts, on which we train Proceed to learn the relationships between concept drifts and desirable parameter adjustments. When confronted with a reoccurring concept drift that has been learned among the synthetic concept drifts, Proceed can effectively tailor the model to the test sample.
@@
  • We propose Proceed, a proactive model adaptation framework for online time series forecasting under concept drift. Proceed first estimates the latent representation of concept drift and then customizes parameter adjustments that adapt the parameters to the concept of the test sample in advance of online prediction.
@@ 2. Definitions and Problem
In online forecasting scenarios, time series data are observed sequentially. Due to the dynamic nature of real-world processes, underlying data distributions are subject to constant change. Consequently, a forecast model trained on historical data may encounter difficulties when confronted with new, evolving patterns. This issue is referred to as the concept drift challenge, which can substantially affect model performance over time (Gama et al., 2014). To mitigate concept drifts, it is crucial to adapt the forecast model continuously to assimilate new concepts presented in the incoming time series. + Formally, the online model adaptation problem is defined as follows.
@@ 3.2. Baseline Variants
  • PatchTST (Nie et al., 2023): PatchTST is one of the state-of-the-art time series forecasters, which models cross-time dependence by Transformer blocks.
@@ 3.3. Evaluation Metrics
where $T_{\text{test}}$ represents all time steps of the test set and $\hat{\mathbf{Y}}_{t}\in\mathbb{R}^{N\times H}$ denotes the predicted result at time $t$. A lower error indicates higher forecasting performance. Each experiment is repeated 3 times with different random seeds and we report the average results. Due to the page limit, we provide the MAE results in Appendix D.1.
@@
Method Dataset $H$ ETTh2 ETTm1 Weather ECL Traffic
@@
  • In Appendix D.1, we report the MAE results in Table 7.
@@ 4.1. Solution Overview
    Feedback-based model adaptation. Given a forecast model $\mathcal{F}$ parameterized by $\boldsymbol{\theta}_{t-H-1}$, we redo forecasting by $\hat{\mathbf{Y}}_{t-H}=\mathcal{F}(\mathbf{X}_{t-H};\boldsymbol{\theta}_{t-H-1})$. Next, we use the forecast error $\|\hat{\mathbf{Y}}_{t-H}-{\mathbf{Y}}_{t-H}\|^{2}_{2}$ to update $\boldsymbol{\theta}_{t-H-1}$ into $\boldsymbol{\theta}_{t-H}$ by gradient descent. The subscript of $\boldsymbol{\theta}$ indicates that the parameters have fit $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$.
@@ 4.3. Proactive Model Adaptation
It is also worth mentioning that our method is applicable to other kinds of model parameters. For a convolution filter $\boldsymbol{\theta}^{(\ell)}\in\mathbb{R}^{d_{\textit{in}}\times d_{\textit{out}}\times d_{\kappa}}$ with $d_{\kappa}$ kernels, we multiply each kernel with ${\boldsymbol{\alpha}_{t}^{(\ell)}}^{\top}\boldsymbol{\beta}_{t}^{(\ell)}$. For a bias term $\boldsymbol{\theta}^{(\ell)}\in\mathbb{R}^{d_{\textit{out}}}$, we only generate $\boldsymbol{\beta}_{t}^{(\ell)}$.
@@ 4.4. Mini-batch Training
Proceed $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$ GD Adjusting all parameters based on $\mathbf{X}_{t-H}$ and $\mathbf{X}_{t}$
@@ 4.5. Discussion
Prior approaches to online time series forecasting make model adaptation passively based on forecasting feedback on previous samples, which mainly focus on effectively learning recent data patterns. In Table 2, we compare existing methods and Proceed in terms of the training data and the model adaptation techniques at each online time $t$. Most methods use one newly acquired training sample, while SOLID (Chen et al., 2024) selects more recent samples that share similar lookback windows with the test sample and are assumed to share similar concept. SOLID and Proceed simply fine-tune the forecast model by gradient descent, while FSNet and OneNet further generates additional parameter adjustments based on the forecasting feedback. Since the feedback is delayed by $H$ steps and cannot reveal the test concept, we propose a novel step called proactive model adaptation which aims to mitigate the effects of concept drift between the training sample and the test sample. As this core step is orthogonal to existing methods, Proceed can also incorporate the data augmentation technique of SOLID and the feedback-based model adaptation techniques of FSNet and OneNet.
@@ 5.1. Experimental Settings
Datasets. As introduced in Sec. 5.1, we use five popular benchmarks , of which more details are provided in Appendix C.1. We conform to the dataset split ratio of FSNet and OneNet, i.e., we split the datasets by 20:5:75 for training, validation, and testing. The rationale is that online learning is of great practical value in scenarios of limited training data, when pretrained forecast models are not adept at handling new concepts during long-term online services.
@@
Online Learning Baselines. We compare our Proceed with OGD (Zinkevich, 2003), a naïve online gradient descent algorithm, and the state-of-the-art online model adaptation methods, including FSNet (Pham et al., 2023), OneNet (Zhang et al., 2023), and SOLID. During the online phase, all methods adopt the Practical strategy described in Sec. 3.2 and update parameters with each new training sample by one-step gradient descent. Among the model adaptation methods, FSNet is specially designed for TCN, while OneNet is based on online ensembling and is model-agnostic. Thus, we implement different variants of OneNet by combining each forecast model with an FSNet. As for SOLID, its official implementation learns linear probing with the pretrained parameters at each update time and does not inherit the fine-tuned parameters from the last update. As the datasets in our experiments have a much longer online phase, we implement a variant called SOLID++ that continually fine-tunes all model parameters to consolidate them with new knowledge across the online data. Empirically, SOLID++ performs better than SOLID in our evaluation setting.
@@ 5.2.1. Effectiveness
To verify the effectiveness of our proposed framework, we compare the forecasting performance of Proceed and other baselines on the five popular datasets. We repeat each experiment with 3 different random seeds and report the average results. As shown in Table 3-4, Proceed achieves the best performance in most cases, reducing the average forecast errors of all models without online learning by 42.3%, 10.3%, 12.9% for TCN, PatchTST, and iTransformer, respectively. Such improvements are considerable, as PatchTST and iTransformer are state-of-the-art time series forecasters. Also, it is worth mentioning that all forecast models have been enhanced by RevIN (Kim et al., 2022), which is another popular approach to concept drift by reducing the distribution shifts w.r.t. the mean and standard deviation of time series via data adaptation. Since our model adaptation method can still achieve remarkable improvements, the results suffice to support our claim that time series forecast models need model adaptation to handle more complex concept drifts.
Figure 5. Efficiency comparison on the Traffic dataset ($H=96$). The horizontal axis is the online forecasting latency (millisecond) between updating the model and obtaining online predictions. We calculate the average latency over 30 runs. The vertical axis is the average MSE on test data. The size of each circle represents the peak amount of GPU memory occupation (GB).

### v2

@@ Abstract.
Time series forecasting always faces the challenge of concept drift, where data distributions evolve over time, leading to a decline in forecast model performance. Existing solutions are based on online learning, which continually organize recent time series observations as new training samples and update model parameters according to the forecasting feedback on recent data. However, they overlook a critical issue: obtaining ground-truth future values of each sample should be delayed until after the forecast horizon. This delay creates a temporal gap between the training samples and the test sample. Our empirical analysis reveals that the gap can introduce concept drift, causing forecast models to adapt to outdated concepts. In this paper, we present Proceed, a novel proactive model adaptation framework for online time series forecasting. Proceed first estimates the concept drift between the recently used training samples and the current test sample. It then employs an adaptation generator to efficiently translate the estimated drift into parameter adjustments, proactively adapting the model to the test sample. To enhance the generalization capability of the framework, Proceed is trained on synthetic diverse concept drifts. Extensive experiments on five real-world datasets across various forecast models demonstrate that Proceed brings more performance improvements than the state-of-the-art online learning methods, significantly facilitating forecast models’ resilience against concept drifts. Code is available at https://github.com/SJTU-DMTai/OnlineTSF.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, Canada.isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data stream miningccs: Computing methodologies Online learning settings
@@ 1. Introduction
Time series forecasting has been a prevalent task in numerous fields such as energy (Boussif et al., 2023), retail (Böse et al., 2017), climate (Bi et al., 2023) and finance (Zhao et al., 2023). Recent years have witnessed a surge of deep learning-based forecast models (Nie et al., 2023; Liu et al., 2024; Zhou et al., 2023; Woo et al., 2024; Han et al., 2024) that take past time series observations to predict values in the next $H$ steps, where $H$ is called forecast horizon. Due to the dynamic nature of the environment, latent concepts that influence observation values (e.g., social interest (Bai et al., 2023), stock market sentiment (Li et al., 2022)) often change over time. This ubiquitous phenomenon is known as concept drift (Gama et al., 2014). In the presence of concept drift, future test data may not follow a similar data distribution as the historical training data, causing degradation in forecasting performance (You et al., 2021; Zhao et al., 2023).
@@
Online learning is commonly used to mitigate the effects of concept drift. The key consideration is to continually transform the newly observed time series into a set of training samples and adjust model parameters by minimizing the forecast errors on the new training samples. In addition to the standard fine-tuning technique, recent works (Pham et al., 2023; Zhang et al., 2023) proposed more advanced model adaptation techniques, which focus on how to effectively adapt to recent data by utilizing forecasting feedback (e.g., errors or gradients) on the new training samples. Among them, FSNet (Pham et al., 2023) monitors the gradients in previous fine-tuning, transforms them into parameter adjustments, and tailors a new forecast model to the current training samples. OneNet (Zhang et al., 2023) is an online ensembling network that generates ensemble weights to combine two forecast models and dynamically adjusts the ensemble weights and the models’ parameter weights according to the forecast errors. However, existing online learning methods overlook the fact that the ground truth of each prediction is not available immediately but is delayed after the forecast horizon.
@@
As illustrated in Fig. 1, an online learning task inherently has a $H$-step feedback delay in time series forecasting, resulting in a temporal gap (at least $H$ steps) between available training samples and the test sample. Formally, at each online time $t$, the current horizon window $\mathbf{Y}_{t}$, which have not been observed yet, has overlap with $\mathbf{Y}_{t-H+1},\dots,\mathbf{Y}_{t-1}$. These samples cannot provide complete forecasting feedback and supervisions for the performant, prevailing time series forecasting models which predict $H$-step values directly (Zeng et al., 2023; Nie et al., 2023; Liu et al., 2024). Hence, the available training samples for online learning at time $t$ is $\mathcal{D}_{t-}=\{(\mathbf{X}_{t^{\prime}},\mathbf{Y}_{t^{\prime}})\mid t^{\prime}\leq t-H\}$, which has a temporal distance to the test sample $(\mathbf{X}_{t},\mathbf{Y}_{t})$. The temporal gap issue suggests that addressing concept drift by fitting the forecast model to the recent training samples may be insufficient, since concept drift may also occur over the $H$-step temporal gap between the recent training samples and the current test sample.
@@
Figure 1. (a) Example of feedback delay when $H=3$. At online time $t=5$, we use observation values $\mathbf{X}_{5}=\{\mathbf{v}_{1},\cdots,\mathbf{v}_{5}\}$ to forecast future values $\mathbf{Y}_{5}=\{\mathbf{v}_{6},\mathbf{v}_{7},\mathbf{v}_{8}\}$, while the ground truth is known until $t=8$ and the feedback of forecasting performance arrives with 3-step delay. (b) A temporal gap always exists between new training samples $\mathcal{D}_{t-}$ and the test sample $\mathbf{X}_{t}$, where concept drift may occur.
@@
To investigate the impact of this temporal gap on forecasting performance, we examine two online learning strategies using real-world time series datasets (detailed in Section 3). The first strategy fine-tunes model parameters using the latest available training sample at time $t$, which is $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$. The second strategy (Pham et al., 2023; Zhang et al., 2023) omits the feedback delay and adapts the model using $(\mathbf{X}_{t-1},\mathbf{Y}_{t-1})$, which is infeasible in practice due to future information leakage. The empirical results show that the average forecast error of the first strategy is approximately double that of the second strategy, indicating that even the latest observed time series pattern differs from the test sample notably. Moreover, the performance gap between the two strategies becomes more significant with a longer horizon, suggesting that the concept drift becomes more pronounced over longer time intervals. These findings highlight the presence of a substantial concept drift between the practical training samples and the test sample, limiting the effectiveness of current online learning techniques that adapt model parameters to the potentially outdated concepts and leave the concept drift unresolved.
@@
[…]
@@
[…]
@@ 2. Definitions and Problem
In online forecasting scenarios, time series data are observed sequentially. Due to the dynamic nature of real-world processes, underlying data distributions are subject to constant change. Consequently, a forecast model trained on historical data may encounter difficulties when confronted with new, evolving patterns. This issue is referred to as the concept drift challenge, which can substantially affect model performance over time (Gama et al., 2014). To mitigate concept drifts, it is crucial to adapt the forecast model continuously to assimilate new concepts presented in the incoming time series. Formally, the online model adaptation problem is defined as follows.
@@ 3.2. Baseline Variants
  • PatchTST (Nie et al., 2023): PatchTST is one of the state-of-the-art time series forecasters, which models temporal patterns by Transformer.
@@ 3.3. Evaluation Metrics
where $T_{\text{test}}$ represents all time steps of the test set and $\hat{\mathbf{Y}}_{t}\in\mathbb{R}^{N\times H}$ denotes the predicted result at time $t$. A lower error indicates higher forecasting performance. Each experiment is repeated 3 times with different random seeds and we report the average results. Due to the page limit, we provide the MAE results in Appendix E.1.
@@
Dataset $H$ ETTh2 ETTm1 Weather ECL Traffic
@@
  • In Appendix E.1, we report the MAE results in Table 7.
@@ 4.1. Solution Overview
    Online Fine-tuning. Given a forecast model $\mathcal{F}$ parameterized by $\boldsymbol{\theta}_{t-H-1}$, we redo forecasting by $\hat{\mathbf{Y}}_{t-H}=\mathcal{F}(\mathbf{X}_{t-H};\boldsymbol{\theta}_{t-H-1})$. Next, we use the forecast error $\|\hat{\mathbf{Y}}_{t-H}-{\mathbf{Y}}_{t-H}\|^{2}_{2}$ to update $\boldsymbol{\theta}_{t-H-1}$ into $\boldsymbol{\theta}_{t-H}$ by gradient descent. The subscript of $\boldsymbol{\theta}$ indicates that the parameters have fit $(\mathbf{X}_{t-H},\mathbf{Y}_{t-H})$.
@@ 4.3. Proactive Model Adaptation
We also apply our method to convolution filters and bias terms. For a convolution filter $\boldsymbol{\theta}^{(\ell)}\in\mathbb{R}^{d_{\textit{in}}\times d_{\textit{out}}\times d_{\kappa}}$ with $d_{\kappa}$ kernels, we multiply each kernel with ${\boldsymbol{\alpha}_{t}^{(\ell)}}^{\top}\boldsymbol{\beta}_{t}^{(\ell)}$. For a bias term $\boldsymbol{\theta}^{(\ell)}\in\mathbb{R}^{d_{\textit{out}}}$, we only generate $\boldsymbol{\beta}_{t}^{(\ell)}$.
@@ 4.4. Mini-batch Training
As our model adapter learns the transition probability $P(\boldsymbol{\theta}_{t}\mid\boldsymbol{\theta}_{s},\mathbf{X}_{s},\mathbf{X}_{t})$, we can follow the proof in the previous work (Bai et al., 2023) to show that it is feasible for $\widehat{\boldsymbol{\theta}}_{t}$ in Eq. (4) to yield lower forecasting error than $\boldsymbol{\theta}_{t-H}$ (see Appendix A).
@@
[…]
@@ 4.5. Discussion
Prior approaches to online time series forecasting make model adaptation passively based on feedbacks in forecasting previous samples, which mainly focus on learning recent data patterns. In Table 2, we compare existing methods and Proceed in terms of the training data and the model adaptation techniques at each online time $t$. Most methods use one newly acquired training sample, while SOLID (Chen et al., 2024) selects more recent samples that share similar lookback windows with the test sample and are assumed to share similar concept. SOLID and Proceed simply fine-tune the forecast model by gradient descent, while FSNet and OneNet further generates additional parameter adjustments based on the forecasting feedback. Since the feedback is delayed by $H$ steps and cannot reveal the test concept, we propose a novel step called proactive model adaptation which aims to mitigate the effects of concept drift between the training sample and the test sample. As this core step is orthogonal to existing methods, Proceed can incorporate the data augmentation technique of SOLID and the feedback-based model adaptation techniques of FSNet and OneNet.
@@ 5.1. Experimental Settings
Datasets. As introduced in Sec. 3.1, we use five popular benchmarks, of which more details are provided in Appendix C.1. We conform to the dataset split ratio of FSNet and OneNet, i.e., we split the datasets by 20:5:75 for training, validation, and testing. The rationale is that online learning is of great practical value in scenarios of limited training data, when pretrained forecast models are inadept at handling new concepts during long-term online service.
@@
Online Learning Baselines. We compare Proceed with GD, a naïve online gradient descent method, and the state-of-the-art online model adaptation methods, including FSNet, OneNet, and SOLID. We reimplement FSNet and OneNet by adopting the Practical strategy described in Sec. 3.2. Furthermore, FSNet is specially designed for TCN, while OneNet is based on online ensembling and is model-agnostic. We implement multiple variants of OneNet by combining each forecast model with an FSNet. As for SOLID, its original work learns linear probing with the pretrained parameters at each update time and does not inherit the fine-tuned parameters from the last update. As our datasets have a much long online phase, we implement a variant called SOLID++ that continually fine-tunes all model parameters across the online data. Empirically, SOLID++ performs better than SOLID in our evaluation setting.
@@ 5.2.1. Effectiveness
To verify the effectiveness of our proposed framework, we compare the predictive performance of Proceed and other baselines on the five popular datasets. We repeat each experiment with 3 runs and report the average results. As shown in Table 3-4, Proceed achieves the best performance in most cases, reducing the average forecast errors of all models without online learning by 42.3%, 10.3%, 12.9% for TCN, PatchTST, and iTransformer, respectively. Also, it is worth mentioning that all forecast models have been enhanced by RevIN (Kim et al., 2022), a data adaptation approach to concept drift by reducing the distribution shifts w.r.t. the mean and standard deviation of time series. Since all online learning methods can still achieve remarkable improvements, it suffices to support our claim that time series forecast models need model adaptation to handle more complex concept drifts.